OpenStudy (anonymous):

David has a hose that is 27 ft long. He wants to cut it into two pieces so that the longer piece is 3 ft more than 2 times the length of the shorter piece. How long will each piece of hose be? Shorter piece: ft Longer piece: ft

OpenStudy (anonymous):

@iGreen @gelow @sleepyhead314 @sleepyjess

OpenStudy (sleepyjess):

First, what would the equation be to solve this?

OpenStudy (sleepyhead314):

make this into an equation 3 more than 2 times the short equals 27 so 3 + 2S = 27 and S + L = 27 (the two pieces added up equal 27) do you know how to use these two equations to solve for the variables? :)

OpenStudy (anonymous):

sorta..

OpenStudy (sleepyhead314):

let's use the 3 + 2S = 27 first can you solve for S ? :3

OpenStudy (anonymous):

um... 2S=27-3?

OpenStudy (sleepyhead314):

good! :) then keep going :D

OpenStudy (anonymous):

2S=24 s=12?

OpenStudy (sleepyhead314):

yes! yay you got it! hehe ^_^ then now that we have s = 12 let's plug it into S + L = 27 to solve for L can you do that? :3

OpenStudy (anonymous):

27-12=15 so is it 12 and 15?

OpenStudy (sleepyhead314):

correct! :D tah dahhhh :)

OpenStudy (anonymous):

thanks Sleepy. :D

OpenStudy (sleepyhead314):

no problem ^_^